OpenStudy (anonymous):

simplify to one natural logarithmic expression: 3ln(a)+2*ln(b)-4ln(c) i need a step by step answer please!!

sam (.sam.):

\[ \log(36)-5\log(4)+\log(20) \] \[ \log(36)-\log(4^5)+\log(20) \] \[ \log(36)-\log(1024)+\log(20) \] \[ \log(\frac{36}{1024} \times 20) \]

sam (.sam.):

Do the division and multiplication , then simplify log